# The $\infty$-S test via quantile affine LASSO

### v2

@@ Abstract
The nonparametric sign test dates back to the early 18th century with a data analysis by John Arbuthnot. It is an alternative to Gosset’s more recent $t$-test for consistent differences between two sets of observations. Fisher’s $F$-test is a generalization of the $t$-test to linear regression and linear null hypotheses. Only the sign test is robust to non-Gaussianity. Gutenbrunner et al. (1993) derived a version of the sign test for linear null hypotheses in the spirit of the F-test, which requires the difficult estimation of the sparsity function. We propose instead a new sign test called $\infty$-S test via the convex analysis of a point estimator that thresholds the estimate towards the null hypothesis of the test.
Keywords: LASSO, nonparametric statistical test, rescaling, robustness, total variation.
@@ 1 Introduction
Given a goodness-of-fit measure such as the log-likelihood, the score test calculates the $\ell_{2}$-norm (weighted by the inverse of Fisher information matrix) of the gradient evaluated at the null hypothesis $H_{0}$ and evaluates whether it is abnormally large compared to a chi-squared distribution. The idea is that if the restricted (under the constraint defined by $H_{0}$) estimator is close the free optimum of the goodness-of-fit measure, the score should not differ from zero significantly. The main advantage of the score test is that it requires only the calculation of the restricted estimate. The score test is equivalent to the Lagrange multiplier test (Silvey, 1959), and to the F-test (Kohler, 1982) for Gaussian linear regression which assumes the data $(X,\boldsymbol{y})$ are related according to
@@
where $X$ is an $n\times p$ matrix and $\boldsymbol{y}\in{\mathbb{R}}^{n}$ is a response, $\boldsymbol{e}$ are errors i.i.d. sampled from a Gaussian distribution ${\rm N}(0,\sigma^{2})$. Given an $m\times p$ matrix $A$ of rank $m$ and a vector $\boldsymbol{b}\in{\mathbb{R}}^{m}$, the F-test considers the linear null and alternative hypotheses
$H_{0}:\ A\boldsymbol{\beta}=\boldsymbol{b}\quad{\rm against}\quad H_{1}:\ A\boldsymbol{\beta}\neq\boldsymbol{b}.$ (2)
We assume that ${\rm rank}(X)=\min(n,p)$ and ${\rm rank}(X)-m<n$, for instance, the situations where $p<n$ or where $p\geq n$ and $m=p$. To derive a pivotal statistic that depends not only on $(X,\boldsymbol{y})$ but also on $(A,\boldsymbol{b})$, Fisher’s F-test uses the Lagrange multiplier technique to define the statistic
$F=\frac{(A\hat{\boldsymbol{\beta}}-\boldsymbol{b})^{\rm T}[A(X^{\rm T}X)^{-1}A^{\rm T}]^{-1}(A\hat{\boldsymbol{\beta}}-\boldsymbol{b})/m}{(\boldsymbol{Y}-X\hat{\boldsymbol{\beta}})^{\rm T}(\boldsymbol{Y}-X\hat{\boldsymbol{\beta}})/(n-p)},$
where $\hat{\boldsymbol{\beta}}=(X^{\rm T}X)^{-1}X^{\rm T}\boldsymbol{Y}$ is the least squares solution of the full model. Under the null $H_{0}$ and for Gaussian noise, the statistic is pivotal with distribution $F_{m,n-p}$. The F-test is not robust to non-Gaussian errors however. Consequently the level of the F-test may not match the nominal level.
Assuming a noise distribution through the choice of a likelihood, the tests discussed above are parametric and may therefore be sensitive to deviation from the assumed distribution. Another class of tests is nonparametric. The oldest nonparametric test (beginning of 18th century) is certainly the sign test employed in the same setting as the paired $t$-test. The advantage of the sign test is that it does not assume Gaussianity, and therefore is robust to deviation from this assumption. Nonparametric tests often make the relaxed assumption that the errors $\boldsymbol{e}$ are i.i.d. sampled from a random variable with zero median. The advantage of nonparametric tests is that they match the nominal level under weak assumptions. The drawback is their lack of power; the signed-rank test is a weighted version of the sign test that gives more power to it, while remaining nonparametric (Wilcoxon, 1945).
Gutenbrunner et al. (1993) extended the sign-based tests for testing equal median of several populations to the linear regression setting (1), where now the errors $\boldsymbol{e}$ are sampled from a random variable with zero median and scale parameter $\xi$ (e.g., the standard deviation for the Gaussian distribution and the median absolute deviation for the Cauchy distribution). Splitting $X=[X_{1},X_{2}]$ and correspondingly $\boldsymbol{\beta}=(\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2})$ with $X_{1}$ and $X_{2}$ have $n$ rows and $p_{1}$ and $p_{2}$ columns, Gutenbrunner et al. (1993) allow to test the exclusion null and alternative hypotheses
$H_{0}:\ \boldsymbol{\beta}_{2}=\boldsymbol{0}\quad{\rm against}\quad H_{1}:\ \boldsymbol{\beta}_{2}\neq\boldsymbol{0}.$ (3)
Letting $O_{d_{1},d_{2}}$ be the zero matrix of dimension $d_{1}\times d_{2}$ and $I_{d_{1}}$ be the identity matrix of dimension $d_{1}\times d_{1}$, these exclusion hypotheses correspond to the general hypotheses (2) with $A=[O_{p_{2}\times p_{1}},I_{p_{2}}]$, $\boldsymbol{b}=\boldsymbol{0}_{p_{2}}$ and $m=p_{2}$. In the spirit of the F-test, they derive a statistic
$T_{n}=S_{n}^{\rm T}Q_{n}^{-1}S_{n}/C_{\varphi}\rightarrow_{d}\chi^{2}_{p_{2}},$ (4)
where $S_{n}$ is the score under $H_{0}$ and $C_{\varphi}$ is some constant; see Gutenbrunner et al. (1993, eq. (2.9)) for details. Their test statistic is the weighted $\ell_{2}$-norm of the score, and the rescaling by $Q_{n}^{-1}$ makes the statistic asymptotically pivotal and $\chi^{2}$ distributed. For that reason we call their test the $\chi^{2}$ sign test. In practice $Q_{n}$ must be estimated which entails estimating the sparsity function $s(\tau)=(f(F^{-1}(\tau)))^{-1}$, where $f$ and $F$ are the unknown density and c.d.f. of the errors, respectively, and $\tau\in(0,1)$ is the quantile of interest; this additional nuisance parameter prevents the test statistic to be close to the asymptotic $\chi^{2}$ distribution even for large sample size $n$ and makes the implementation unstable, as we observe in our empirical experiments in Section 5.
Interestingly the exclusion hypotheses (3) allow to test the general hypotheses (2) for any $(A,\boldsymbol{b})$ using the following transformation. Letting $\tilde{A}$ be a $(p-m)\times p$ matrix that completes $A$ such that $[\tilde{A}^{\rm T},A^{\rm T}]$ is nonsingular and $A\tilde{A}^{\rm T}=O_{m\times(p-m)}$, they transform the data to $\tilde{\boldsymbol{y}}=\boldsymbol{y}-XA^{\rm T}(AA^{\rm T})^{-1}\boldsymbol{b}$, $\tilde{X}_{1}=X\tilde{A}^{\rm T}(\tilde{A}\tilde{A}^{\rm T})^{-1}$ and $\tilde{X}_{2}=XA^{\rm T}(AA^{\rm T})^{-1}$ and apply their test with $(\tilde{X}_{1},\tilde{X}_{2},\tilde{\boldsymbol{y}})$ in place of $(X_{1},X_{2},\boldsymbol{y})$ to test (2). Mathematically elegant, this transformation is not innocuous because computationally cumbersome. Indeed this transformation requires to find a matrix $\tilde{A}$ with the desired properties and invert $\tilde{A}\tilde{A}^{\rm T}$. Take the example where one wants to test $\beta_{1}=\beta_{2}$: $A$ is a $1\times p$ matrix, so $\tilde{A}$ is $(p-1)\times p$ and $\tilde{A}\tilde{A}^{\rm T}$ is $(p-1)\times(p-1)$, which may be difficult to invert with today’s data where $n$ and $p$ are large. We see that this indirect way for handling the general linear hypotheses is tedious.
Like the F-test, we derive instead a statistic that directly tackles the desired null hypotheses (2), but as opposed to the F-test, we do not seek an asymptotic $\chi^{2}$ distribution. Our derivation of the $\infty$-S test is based on the least absolute deviation version of the affine LASSO point estimator (Sardy et al., 2022). Affine LASSO is a regularized estimator which regularization parameter can be seen as a Lagrange multiplier to constrain the null hypothesis. So the $\infty$-S test is both a score test and a Lagrange multiplier test.
The article is organized as follows. In Section 2, we derive the $\infty$-S test by defining the affine LAD LASSO (least absolute deviation least absolute shrinkage and selection operator), deriving its zero-thresholding function, and defining the sign-based test statistic that is asymptotically pivotal with respect the nuisance parameters. In Section 3, we discuss optimization issues related to LAD regression with linear constraints and affine LAD LASSO. In Section 4, we extend the $\infty$-S test to testing a quantile of the distribution through the definition of the quantile affine LASSO estimator. In Section 5, we apply the $\infty$-S test to four settings: two simple settings to retrieve existing sign tests, a quantile total variation sign test to test for a jump in the tail of time series, and an empirical robustness analysis to non-Gaussianity of the $\infty$-S test and the F-test to match the desired nominal level of the test. The proofs are postponed to an Appendix.
2 The $\infty$-S test
The affine LAD-LASSO estimator is a solution to
$\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}.$ (5)
@@ 2 The $\infty$-S test
The larger $\lambda$, the more this estimator thresholds towards zero the entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ corresponding to the null hypothesis. As stated in the following theorem, there is an interval $[\lambda_{0},\infty)$ for which the solution to (5) satisfies the condition of $H_{0}$ for any $\lambda\in[\lambda_{0},\infty)$. The smallest such $\lambda$ is a function of the data and has a closed form expression called the zero-thresholding function.
@@
$\hat{\boldsymbol{\beta}}_{H_{0}}\in\arg\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}\quad{\rm s.t.}\quad A\boldsymbol{\beta}=\boldsymbol{b}.$ (6)
@@
where ${\boldsymbol{\omega}}$ is the dual variable associated to (5), which are the sign of $X\hat{\boldsymbol{\beta}}_{H_{0}}-{\boldsymbol{y}}$ of the constrained LAD (6).
Since satisfying the conditions of $H_{0}$ are equivalent to $\|A\boldsymbol{\beta}-\boldsymbol{b}\|=0$ for any norm, the zero-thresholding function (7) is the (scalar) Lagrange multiplier to satisfy the constraint $\|A\boldsymbol{\beta}-\boldsymbol{b}\|=0$; using the $\ell_{1}$-norm, the Lagrange multiplier that achieves this is finite. Instead of the $\ell_{1}$-norm, we could have used the $\ell_{2}$-norm corresponding to affine LAD-group LASSO (Yuan and Lin, 2006); the dual of the $\ell_{2}$-norm being the $\ell_{2}$-norm, the corresponding zero-thresholding function would have been $\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{2}$. More generally, one can use the $\ell_{q}$-norm, which corresponding zero-thresholding function is $\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{q/(q-1)}$.
The zero-thresholding function leads to the statistic of the $\infty$-S test, which is asymptotically pivotal, as stated in the following theorem.
@@
Assuming the LAD coefficient estimates are asymptotically Gaussian centered around the true coefficients, the test statistic $S=\lambda_{0}(\boldsymbol{Y},X,A)$, where $\boldsymbol{Y}$ is the response random vector under $H_{0}$ is asymptotically pivotal.
The test statistic $S$ involves the score of the LAD loss under $H_{0}$ like the statistic $T_{n}$ of Gutenbrunner et al. (1993) with two noticable differences: the test statistic $S$ uses the infinite norm instead of the weighted $\ell_{2}$-norm, and its distribution does not require estimating the sparsity function of Gutenbrunner et al. (1993). The sparisty function is only needed to make the statistic asymptotically $\chi^{2}$under $H_{0}$, not to make the statistic pivotal. Indeed, as Theorem 2 shows, the score is already asymptotically pivotal. There is an important change of paradigm between the generalization of the sign test by Gutenbrunner et al. (1993) and our new $\infty$-S test: their test has an asymptotic listed distribution (e.g., $F$ or $\chi^{2}$), while our test has an unlisted distribution. Nevertheless, with today’s computer, one can easily sample the distribution of the test statistic $S=\lambda_{0}(\boldsymbol{Y},X,A)$ by Monte Carlo and estimate the test critical value as the empirical $(1-\alpha)$-quantile of $S$. The p-value is also estimated as the percentage of Monte Carlo simulated random variables larger than $\lambda_{0}(\boldsymbol{y},X,A)$.
The asymptotic Gaussianity of the LAD coefficient estimates required by Theorem 2 has been proved by Pollard (1991), so we can define the $\infty$-S test.
@@
Consider the linear model (1). To test (2) at a prescribed level $\alpha\in(0,1)$, let $c_{\alpha}=F_{S}^{-1}(1-\alpha)$, where $S=\lambda_{0}(\boldsymbol{Y},X,A)$ and $\boldsymbol{Y}$ is the response random vector for the linear model under $H_{0}$ and the function $\lambda_{0}$ is given in (7). The $\infty$-S test function is defined by
$\phi({\bf y},X,A)=\left\{\begin{array}[]{rl}0&\lambda_{0}(\boldsymbol{y},X,A)\leq c_{\alpha}\\ 1&{\rm otherwise}\end{array}\right.,$
where ${\boldsymbol{y}}$ is the data response.
Based on Theorem 2, the level of the $\infty$-S test matches asymptotically the nominal level $\alpha$. The $\infty$-S test looks at the sign of the residuals, not their amplitude, causing some lack of power of the $\infty$-S test with respect to the F-test for Gaussian models. Weighting the sign $\boldsymbol{\omega}$ of the residuals by the rank of the absolute value of the respective residuals leads to the $\infty$-rankS test that has more power.
To completely define the test, the rescaling of $(A,\boldsymbol{b})$ must be performed to achieve homogeneous power, that it, to not favor certain alternative hypotheses to the detriment of others. Indeed, the matrices $A$ and $X$ create heteroschedasticity in the random vector $\boldsymbol{W}$ in the test statistics $S=\|\boldsymbol{W}\|_{\infty}$ with $\boldsymbol{W}=|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}|$ before taking the maximum of all entries. Consequently, under $H_{1}$, the entries of $\boldsymbol{W}$ with small deviations will be hidden behind the entries with large deviations. To be democratic between potential entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ not being equal to zero, we propose the following rescaling.
@@
Consider the $\infty$-S test with test statistic $S=\|\boldsymbol{W}\|_{\infty}$, where $\boldsymbol{W}=|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}|$ and ${\boldsymbol{\omega}}$ are the sign of the residuals of the constrained LAD under $H_{0}$. Let $F_{W_{k}}$ be the marginal cumulative distribution function of the $k$th entry of $\boldsymbol{W}$ for $k=1,\ldots,m$. Letting $D_{(A,X,{\bf b},\alpha)}$ be the diagonal matrix with $k$th diagonal entry equal to $d_{k}:=F_{W_{k}}^{-1}(1-\alpha)$, homopower rescaling rescales $A$ to $D_{(X,A,{\bf b},\alpha)}A$ and $\boldsymbol{b}$ to $D_{(X,A,{\bf b},\alpha)}\boldsymbol{b}$ before applying the test.
3 Optimization
The cost function of the LAD optimization
$\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}$ (8)
is convex, and it is well known that the optimization can be transformed into a linear program (Koenker and Bassett, 1978). The $\infty$-S test requires solving the constrained LAD optimization (6) that can be transformed into an unconstrained LAD optimization of the form (8).

### v3

@@ Abstract
A novel test in the linear $\ell_{1}$ (LAD) and quantile regressions is proposed, based on the scores provided by the dual variables (signs) arising in the calculation of the (so-called) affine-lasso estimate—a Rao-type, Lagrange multiplier test using the thresholding, towards the null hypothesis of the test, function of the latter estimate.
keywords:
LAD regression, quantile regression, LASSO, Nonparametric statistical test, Rescaling, Robustness, Total variation.
@@ 1 Introduction
We propose a new statistical test of a linear hypothesis in the linear $\ell_{1}$ (LAD) and quantile regression. In that context, the inference is considerably nonparametric, as the null hypothesis in testing is as a rule expressed via certain restriction on the median or pertinent quantile. For instance, the well-known sign test, going back to John Arbuthnot in 1710 (Spr89; Con99), tests the null hypothesis that the common median of independent observations $Y_{i}$ is $\mu$; the test statistic is then the number of positive signs among the signs of all $Y_{i}-\mu$ (in practice, one would like to avoid zeros of $Y_{i}-\mu$, but that is a minor complication that can be dealt with). Apart from the independence assumption, the distributional specification is quite loose; the $Y_{i}$’s even do not have to have the same distribution. Thus, there is a lot of peace of mind in the application of the test, as one has not to put too much faith in overly detailed assumptions. In the traditional terminology, the sign test is dubbed nonparametric, meaning that it is guaranteed to maintain the nominal level under fairly weak assumptions—and at the same time exhibits reasonable power against numerous alternatives; of course, here being nothing for free, the potential cost of such generality may be the loss of efficiency compared to certain parametric tests derived for exactly specified distributions—but only rather in case when the data really follow those.
It is well-known now that the use of the $\ell_{1}$ loss (cost) function in regression predated the introduction of the $\ell_{2}$ one by Gauss and Legendre by about half a century. The computation feasibility of the latter, combined with the flexibility and plausibility of the distributional assumption of normality, eclipsed the early 1760 invention of Boscovich for the centuries to come. However, the computational breakthroughs in linear programming, and among all, the possibility of broadening of the scope of the $\ell_{1}$ approach of the LAD (“Least Absolute Deviations”) regression to quantile regression, the study of conditional quantiles championed by KB78, brought the pendulum of the attention considerably back.
This gained an additional strong momentum when the $\ell_{1}$ loss function emerged also in the penalized formulations of LASSO (Least Absolute Shrinkage and Selection Operator, also known as Basis Pursuit)(CDS99; Tibs:regr:1996). While the original formulations retained the $\ell_{2}$ loss in the lack-of-fit term, to maintain a bond with the prevailing methodology, an attractive synthesis arose in the quantile regression LASSO, its particular median case known as LAD-LASSO (CIS-215377); the latter framework carries a conceptual advantage that both lack-of-fit term and penalty term feature the same form of the loss functions—so that the LAD-LASSO can be written in an augmented $\ell_{1}$ regression form.
All that said, the problem of inference in the $\ell_{1}$ regression is indisputably appealing. The large-sample approach dates back to kolmog31, who already established the asymptotic normality of the sample median for independent, identically distributed observations; this was extended to the regression case by LADasym91, and refined by others. The main hurdle of the implementation of this approach is that the asymptotic variance depends on the reciprocal of the density (that is, “sparsity”) at the (potentially unknown) median of the distribution of the observations (the distribution of the errors in the regression case). The large-sample Wald-type test is then bound to estimate this unknown nuisance parameter; while this is not an impossible task—various strategies proposed in this respect are reviewed in the book of Koenker—it may still constitute a hurdle one would like to bypass.
And indeed: the Rao’s “score” approach to testing (also known as Lagrange multiplier test), elaborated for inference in the $\ell_{1}$/quantile regression by Roger93 via so-called regression rank scores (gutejure92), the dual variables of the primal convex optimization task defining the estimates, brought far-reaching generalizations of the traditional one-sample and two-sample rank tests (including, in particular, the sign test). While it is a general wisdom that these methods avoid the estimation of sparsity—as underlined explicitly by Koenkerin the context of confidence intervals—he2023 mention computational instabilities for large samples and various other problems, and also the estimation of asymptotic variance which could allegedly mar the rank test approach—quoting, however, rather the more abstract account of gutejure92 instead of the more focused Roger93 .
The problem here is pretty much in the eye of beholder, however. Unlike the simple hit-and-miss paradigm of the sign-test, rank tests allow for various flavors depending on what kind of rank scores are adopted; the scores make Wilcoxon tests different from the van der Waerden ones, for example. A human investigator may prefer one to another for the one- and two-sample problems, and accordingly may opt for the corresponding scores also in the $\ell_{1}$/quantile setting. For instance, van der Waerden scores may be preferred by those who believe that the reality is mostly normal (“Gaussian”), apart from occasional erratic disturbances. Thus, once the flavor of the rank test is chosen a priori, the procedure “does not require any nuisance parameter depending on the error distribution to be estimated” (Roger93).
However, when the rank testing is to be performed in an automatic way, the rank scores may have to be elucidated in an “adaptive way”, using the existing theory of optimality of certain scores for certain distributions (for example, as mentioned above, van der Waerden scores work best for the normal distribution) to estimate the right scores directly from the data. Ascending this next step on the ladder brings a need of distributional estimation, potentially including that of sparsity, back into the game.
The desire for automatic, human-free procedures may thus favor a simple solution without a need of additional tuning choices; if such a solution is available, and exhibits decent power compared to the already existing options, it possesses a definite raison-d’être, the aspect we aim at demonstrating below. Section 2 gives the derivation of $\infty$-S test by defining the affine LAD LASSO, elucidating its zero-thresholding function, and subsequently proposing the sign-based test statistic that is asymptotically pivotal with respect to the nuisance parameters. In Section 3, we discuss optimization issues related to LAD regression with linear constraints and affine LAD LASSO; Section 4 extends the $\infty$-S test to testing a quantile of the distribution through the definition of the quantile affine LASSO estimator. In Section 5, the $\infty$-S test is applied in four settings: two simple settings to retrieve existing sign tests, a quantile total variation sign test to test for a jump in the tail of time series, and an empirical robustness analysis to non-Gaussianity of the $\infty$-S test and the F-test to match the desired nominal level of the test. The proofs are postponed to an Appendix.
2 The $\infty$-S test
Consider a linear model
@@ 2 The $\infty$-S test
where $X$ is an $n\times p$ matrix and $\boldsymbol{\beta}$ is a $p\times 1$ vector of unknown parameters. Given an $m\times p$ matrix $A$ of rank $m$ and a vector $\boldsymbol{b}\in{\mathbb{R}}^{m}$, we consider testing
$H_{0}:\ A\boldsymbol{\beta}=\boldsymbol{b}\quad\text{against}\quad H_{1}:\ A\boldsymbol{\beta}\neq\boldsymbol{b}.$ (2)
A new testing procedure, which we propose to call an $\infty$-S test, is based on the least absolute deviation version of the affine LASSO point estimator (SARDY2022107507), a regularized estimator that can be defined in the constrained form as a solution of the problem
$\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}\to\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}!\qquad\text{subject to}\qquad\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}\leq\Lambda.$ (3)
The “Lagrangian” form of the definition used by SARDY2022107507 proclaims the same estimator a solution to
$\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}\to\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}!$ (4)
@@
for a suitable Lagrange multiplier $\lambda$ pertaining to the constraint in (3); the larger $\lambda$, the more this estimator thresholds towards zero the entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ corresponding to the null hypothesis $H_{0}$ in (2). That is, there is an interval $\lambda_{0},\infty)$ for which the solution to (4) satisfies the condition of $H_{0}$ for any $\lambda\in[\lambda_{0},\infty)$. The smallest such $\lambda$ is a function of the data and has a closed form expression which we call the zero-thresholding function.
@@
$\hat{\boldsymbol{\beta}}_{H_{0}}\in\arg\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}\quad\text{\rm subject to}\quad A\boldsymbol{\beta}=\boldsymbol{b}.$ (5)
@@
where $\boldsymbol{\omega}$} is the dual variable associated to (4), the sign of $X\hat{\boldsymbol{\beta}}_{H_{0}}-{\boldsymbol{y}}$ of the constrained LAD (5) whenever $X\hat{\boldsymbol{\beta}}_{H_{0}}-{\boldsymbol{y}}\neq 0$.
As satisfying $H_{0}$ is equivalent to $\|A\boldsymbol{\beta}-\boldsymbol{b}\|=0$ for any norm, the result of the zero-thresholding function (6) is the the constraint $\|A\boldsymbol{\beta}-\boldsymbol{b}\|=0$; with the $\ell_{1}$-norm, the pertinent Lagrange multiplier is finite. Instead of the $\ell_{1}$-norm, we could have used the $\ell_{2}$-norm corresponding to affine LAD-group LASSO (Yuan:Lin:mode:2006); the dual of the $\ell_{2}$-norm being the $\ell_{2}$-norm itself, the corresponding zero-thresholding function would have been $\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{2}$. More generally, one can use the $\ell_{q}$-norm, which corresponding zero-thresholding function is $\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{q/(q-1)}$.
The zero-thresholding function yields the test statistic of the $\infty$-S test, which the following theorem shows is asymptotically pivotal. The $\infty$-S test can be thus considered “Lagrange multiplier test”—that is, the Rao score test— akin to the rank tests developed by Roger93; but, unlike the statistic $T_{n}$ of Roger93, the $\infty$-S test does not require the integration of scores for dual variables coming from different quantile regressions, and avoids also the need to select the rank scores (and subsequent potential estimation of underlying density characterizations to achieve the optimality of those).
@@
Let $\boldsymbol{Y}$ be the response random vector under $H_{0}$. Assuming the LAD coefficient estimates are asymptotically Gaussian centered around the true coefficients, the test statistic $S=\lambda_{0}(\boldsymbol{Y},X,A)$ is asymptotically pivotal.
The asymptotic normality of the LAD coefficient estimates required by Theorem 2 is a well known fact—proved by LADasym91, and refined by others. We are therefore in the position to define the $\infty$-S test now.
@@
The $\infty$-S test function, to test (2) in a linear model  (1) at a prescribed level $\alpha\in(0,1)$, is defined to be
$\phi({\boldsymbol{y}},X,A)=\left\{\begin{array}[]{rl}0&\lambda_{0}(\boldsymbol{y},X,A)\leq c_{\alpha},\\ 1&\text{\rm otherwise},\end{array}\right$.,
where $c_{\alpha}$ is selected so that the test has nominal level $\alpha$.
Based on Theorem 2, the level of the $\infty$-S test matches asymptotically the nominal level $\alpha$. The test itself is being based on the dual variables for the $\ell_{1}$ minimization problem; these variables attain values in [1,1][-1,1] and amount to the sign of residuals if those are not equal to zero. The insensitivity to the magnitude of the residuals makes the test resistant to outlying observations; nonetheless, in model-behaved situations, when the error distribution is exactly normal, a more powerful $\infty$-rankS test can be obtained by weighting the sign $\boldsymbol{\omega}$ of the residuals by the rank of the absolute value of the respective residuals.
In the applications, the sought p-values (or critical values $c_{\alpha}$, if desired) are then easily obtained via Monte Carlo sampling of $S=\lambda_{0}(\boldsymbol{Y},X,A)$, the technology which nowadays provides satisfactory results even in the absence (so far) of a knowledge of the distribution of test statistics in closed form (related, say, to some traditional distributions).
From the practical point of view, we also recommend, in the spirit of SardyISI08, to perform preliminary rescaling of $(A,\boldsymbol{b})$, in order to achieve homogeneous power—that is, the power not favoring certain alternative hypotheses to the detriment of the other ones. Indeed, the matrices $A$ and $X$ create heteroskedasticity in the random vector $\boldsymbol{W}$ in the test statistics $S=\|\boldsymbol{W}\|_{\infty}$ with $\boldsymbol{W}=|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}|$ before taking the maximum of all entries. Consequently, under $H_{1}$, the components of $\boldsymbol{W}$ with small deviations will be hidden behind the components with large deviations. To treat equally potential components of $A\boldsymbol{\beta}-\boldsymbol{b}$ not being equal to zero, we propose the following.
@@
Consider the $\infty$-S test with test statistic $S=\|\boldsymbol{W}\|_{\infty}$, where $\boldsymbol{W}=|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}|$ and $\boldsymbol{\omega}$} are the sign of the residuals of the constrained LAD under $H_{0}$. Let $F_{W_{k}}$ be the marginal cumulative distribution function of the $k$th component of $\boldsymbol{W}$ for $k=1,\ldots,m$. Letting $D_{(A,X,{\bf b},\alpha)}$ be the diagonal matrix with $k$th diagonal element equal to $d_{k}:=F_{W_{k}}^{-1}(1-\alpha)$, the homopower rescaling rescales $A$ to $D_{(X,A,{\bf b},\alpha)}A$ and $\boldsymbol{b}$ to $D_{(X,A,{\bf b},\alpha)}\boldsymbol{b}$ before applying the test.
3 Optimization aspects
The objective function of the LAD optimization task
$\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}\to\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}!$ (7)
is convex, and it is well known that the optimization can be transformed into a linear program (KB78). The $\infty$-S test requires solving the constrained LAD optimization (5) that can be transformed into an unconstrained LAD optimization of the form (7).